# Nonsingular black holes and spherically symmetric objects in nonlinear electrodynamics with a scalar field

### v2

@@ V Theories with ${\cal L}=X+\mu(\phi)F^{n}$
In theories where the Lagrangian ${\cal L}$ contains the nonlinear dependence of $F$, ${\cal L}_{,FF}$ does not vanish in Eq. (68) or Eq. (86). As we showed in Sec. IV, this results in the negative values of $c_{\Omega,V}^{2}$ or $c_{\Omega,\delta A}^{2}$ around $r=0$. If we consider theories in which ${\cal L}$ contains only a linear term in $F$, it is possible to avoid the problem of angular Laplacian instabilities. We also note that the nonlinear dependence of $X$ in ${\cal L}$ leads to the deviation of $c_{r,\delta\phi}^{2}$ from 1. The linear term in $X$ without a direct coupling to $F$ results in the value $c_{r,\delta\phi}^{2}=1$ for both electric and magnetic cases.
@@ V.1 Electric case
$-\frac{1}{2}h\phi^{\prime 2}+\mu(\phi)\left(\frac{A_{0}^{\prime 2}}{2N}\right)^{n}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)+\frac{q_{E}A_{0}^{\prime}}{\sqrt{N}r^{2}}\,.$ (108)
@@
$h(r)$ $=$ $\frac{2\int_{0}^{r}r_{1}^{-\frac{4(n-1)}{2n-1}}\sqrt{N(r_{1})}\bigl{(}\int_{0}^{r_{1}}\sqrt{N(r_{2})}{\rm d}r_{2}\bigr{)}{\rm d}r_{1}}{(2n-1)r^{\frac{2}{2n-1}}N(r)}$ (111)
@@
In general, the set of solutions corresponding to this $\mu(\phi)$, if it is neither empty nor a singleton, will be distinguished by their respective values of mass and charge of the object. [^2] This would mean that, for a fixed $\mu(\phi)$, there would be more families of profiles, e.g., for $N(r)$ and $\phi(r)$. The variation in these parameters may result in different forms of $N(r)$, [^3] Notice, however, that for the RN solution in GR, although the mass and charge vary, the function $N$ is unity over the whole manifold. For perfect fluids, we instead have $N\neq 1$, when a non-negligible pressure is present. which must satisfy the boundary conditions $N^{\prime}(r\to 0)\propto r^{3}$ at the origin and $N\to 1$ at infinity. [^4] In addition, we require the scalar-field dependence $\phi^{\prime}(r\to 0)\propto r$ as well as all the other boundary conditions that define regular objects with asymptotically flat spacetime. Exploring this second avenue, checking the properties of solutions for a fixed $\mu(\phi)$ is an interesting prospect. Still, we believe that its investigation lies beyond the scope of this work and should be addressed in a future study.
@@
where $G$ is the gravitational constant. From Eqs. (127) and (128), for $n\geq 3/2$, the quantity $r(1-h)$ increases at large distances. In this case, we do not have compact SSS objects. For $n$ in the range
@@
the metric component behaves as Eq. (126) and hence $M(r)$ approaches a constant value $-c_{1}/(2G)$. This is the case in which nonsingular SSS objects satisfy the condition for compactness.
@@
where the term proportional to $r^{-{\frac{2}{2n-1}}}$ dominates over $r^{-4}$ for $23/4<n<3/$ 2. As we will see later in Fig. 5, the allowed values of $n$ that are consistent with no-ghost conditions are indeed larger than $3/4$.
@@
In Fig. 2, we plot the ADM mass (129) for seven different values of $n$, with $N_{0}=0.2$. The analytic estimation (126) of $h(r)$ shows that $M$ asymptotically approaches constants for 1/$21/2<n<3/$ 2, whose property can be confirmed in Fig. 2. For $n=3/2$, the ADM mass exhibits the logarithmic growth $M\propto\ln r$ due to the property of Eq. (127). For $n>3/2$, Eq. (128) gives the analytic dependence $M\propto r^{\frac{2n-3}{2n-1}}$. In Fig. 2, we can confirm this power-law growth of $M$ for $n>3/2$ at large distances. Thus, so long as 1/$21/2<n<3/$ 2, the electric SSS configuration without the apparent horizon can be interpreted as a compact object.
@@
Figure 5: The allowed region of parameter space for electric compact objects, represented by the light-blue shaded area. The lower boundary corresponds to the values of $n_{\rm min}$, below which the no-ghost condition is violated for a given $N_{0}$ in the range $10<N_{0}<$ 1. The upper boundary, represented by $n=3/2$, marks the limit beyond which the solutions no longer describe compact objects. Thus, the allowed range for $n$ is $n_{\rm min}\leq n<3/2$. The curve of $n_{\rm min}(N_{0})$ is obtained by interpolating over a set of 99 numerically computed values of $n_{\rm min}$ for $N_{0}$ in the range $0.01\leq N_{0}\leq 0.99$.
@@
The metric component $f=Nh$ differs from $h$ by the factor $N$, whose difference is largest at $r=0$ by the factor $N_{0}=0.2$. Since $N$ grows from $N_{0}$ to 1 for increasing $r$, $f$ approaches $h$ at large distances and they finally converge to the asymptotic values $f\to 1$ and $h\to 1$. The analytic estimation (124) shows that $h(r)$ has the large-distance behavior $h(r)=1+c_{1}/r+c_{2}/r^{2}$ for $n=1$. From Fig. 3, we find that $h(r)<1$ and hence $c_{1}$ is negative. As we see in Fig. 2 for $n=1$, the ADM mass $M$ approaches a constant positive value $-c_{1}/(2G)$ at spatial infinity.
@@
Since we are now considering the case 1/$21/2<n<3/$ 2 with $F>0$, the no-ghost condition for electric compact objects is equivalent to $\mu>0$, i.e.,
$\mu$ $=$ $\frac{2^{n-1}M_{\rm Pl}^{2}[2N(1-h-rh^{\prime})-rhN^{\prime}]}{(2n-1)r^{2}N}$ (139)
×$\times\left[\frac{q_{E}^{2}(2n-1)^{2}N^{2}}{M_{\rm Pl}^{4}n^{2}\{2(rh^{\prime}+h-1)N+rhN^{\prime}\}^{2}}\right]^{n}>0\,.$
For given values of $N_{0}$, we study how the no-ghost condition $\mu>0$ depends on the parameter $n$ within the allowed range 1/$21/2<n<3/$ 2. Our numerical analysis reveals the existence of a minimum value of $n$ for each $N_{0}$, denoted as $n_{\rm min}$, above which $\mu>0$ for all $r$. Conversely, if $n<n_{\rm min}$, $\mu$ becomes negative at some finite distance $r$, so these values of $n$ are excluded.
In Fig. 4, we plot $\mu$ versus $r/r_{0}$ with the choice $N_{0}=0.2$ for five different values of $n$. In this case, if $n<0.887$, the coupling enters the region $\mu<0$ for some values of $r$. For $n>n_{\rm min}=0.887$, we observe in Fig. 4 that $\mu>0$ at any distance $r$. For $N_{0}$ in the range $10<N_{0}<$ 1, we numerically obtain the minimum values of $n$ and plot $n_{\rm min}$ versus $N_{0}$ in Fig. 5. We find that $n_{\rm min}$ mildly increases as a function of $N_{0}$ and approaches 1 for $N_{0}\to 1$. Note that $n_{\rm min}$ is always larger than $1/2$ for any values of $N_{0}$. So long as $n$ is in the range
@@ V.2 Magnetic case
where we have chosen the branch $\phi^{\prime}>0$. The Lagrangian ${\cal L}$ obeys the following relation
$-\frac{M_{\rm Pl}^{2}hN^{\prime}}{2rN}+\mu(\phi)\left(-\frac{q_{M}^{2}}{2r^{4}}\right)^{n}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,.$ (143)
If $n$ is an integer, the second term in Eq. (143) reduces to $[q_{M}^{2}/(2r^{4})]^{n}$ for even $n$ and $-[q_{M}^{2}/(2r^{4})]^{n}$ for odd $n$. If $n$ is not an integer, we should think of $F^{n}$ in the Lagrangian as $(F^{2})^{m}$, i.e., $n=2m$, so that Eq. (143) reduces to
@@
For both integer and non-integer values of $n$, the coupling $\mu(\phi)$ is known in terms of $h$, $N$, and their $r$ derivatives.
@@
The integrated solution to Eq. (146) can be obtained by replacing the power $n$ to $n/(2n-1)$ in Eq. (111), so that
@@
Figure 7: This plot depicts the allowed region of parameter space, indicated by the light-blue shaded area. The upper boundary corresponds to the value of $n_{\rm max}$, beyond which the no-ghost condition is violated for a given $N_{0}$ within the allowed range $10<N_{0}<$ 1. The lower boundary, at $n=3/4$, defines the limit below which the solutions no longer describe a compact object. Consequently, the permissible range for $n$ is $3/4<n\leq n_{\rm max}$.
@@
The metric component $h$ behaves as Eq. (118) around $r=0$, where $n$ should be replaced with $n/(2n-1)$. At large distances, we require that the ADM mass $M$ is converging, whose condition corresponds to $n<3/2$ for the electric configuration. Using the electric-magnetic duality, the condition for compactness of magnetic SSS objects is now given by $n>3/4$. For the choice (120), the large-distance solutions to $h(r)$ are given by the dual of Eq. (124), that is
@@ VI.1 Electric case
around $r=0$. The reason why the leading-order terms of $c_{r,\delta\phi}^{2}$ and $c_{\Omega,V}^{2}$ are negative is attributed to the property $\kappa\propto F^{-1}$. Indeed, we can obtain the value $-1/3$ by substituting the relation $\kappa\propto F^{-1}$ into Eqs. (158) and (159). Even though the dependence of $\kappa$ on $F$ is different for the distance $r$ away from the origin, the Laplacian instability around $r=0$ is sufficient to exclude the above horizonless solution as a stable SSS configuration.
@@ VI.2 Magnetic case
whose leading-order term is negative. Hence the nonsingular SSS object without the horizon is excluded by angular Laplacian instability at large distances. We note that, even without imposing the condition $\tilde{N}_{0}=1$ for asymptotic flatness, the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is $-1/2$. The above results show that, as $r$ increases from 0, $c_{\Omega,\delta A}^{2}$ enters the regime $c_{\Omega,\delta A}^{2}<0$ to approach the asymptotic value $-1/2$. Indeed, we have numerically confirmed this property [^6] This behavior of $c_{\Omega,\delta A}^{2}$ is similar to what happens for hairy BHs present in cubic vector Galileon theories [64]. by choosing $N(r)$ of the form (120).
@@ VII Conclusions
In Sec. IV, we applied the linear stability conditions to NED with a k-essence scalar field. For nonsingular electric SSS objects, we showed that the even-parity vector-field perturbation $V$ is subject to angular Laplacian instability around $r=0$ due to the negative leading term of $c_{\Omega,V}^{2}$. For the nonsingular magnetic configuration, the same angular Laplacian instability arises for the odd-parity vector-field perturbation $\delta A$. Since the fields $V$ and $\delta A$ are coupled to the even-parity gravitational perturbation $\psi$, the background metric regular at $r=0$ cannot remain in a steady state. This means that, in theories given by the Lagrangian ~$\tilde{\cal L}(F)+K(\phi,X)$, the nonsingular SSS objects (including regular BHs) do not exist as stable configurations.

### v3

@@ V Theories with ${\cal L}=X+\mu(\phi)F^{n}$
[…]
@@ V.1 Electric case
$\frac{1}{2}h\phi^{\prime 2}+\mu(\phi)\left(\frac{A_{0}^{\prime 2}}{2N}\right)^{n}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)+\frac{q_{E}A_{0}^{\prime}}{\sqrt{N}r^{2}}\,.$ (108)
@@
$\hskip-28.45274pth(r)$ $=$ $\frac{2\int_{0}^{r}r_{1}^{-\frac{4(n-1)}{2n-1}}\sqrt{N(r_{1})}\bigl(\int_{0}^{r_{1}}\sqrt{N(r_{2})}{\rm d}r_{2}\bigr){\rm d}r_{1}}{(2n-1)r^{\frac{2}{2n-1}}N(r)}$ (111)
@@
In general, the set of solutions corresponding to this $\mu(\phi)$, if it is neither empty nor a singleton, will be distinguished by their respective values of mass and charge of the object. [^3] This would mean that, for a fixed $\mu(\phi)$, there would be more families of profiles, e.g., for $N(r)$ and $\phi(r)$. The variation in these parameters may result in different forms of N(r)N(r), [^4] Notice, however, that for the RN solution in GR, although the mass and charge vary, the function $N$ is unity over the whole manifold. For perfect fluids, we instead have $N\neq 1$, when a non-negligible pressure is present. which must satisfy the boundary conditions $N^{\prime}(r\to 0)\propto r^{3}$ at the origin and $N\to 1$ at infinity. [^5] In addition, we require the scalar-field dependence $\phi^{\prime}(r\to 0)\propto r$ as well as all the other boundary conditions that define regular objects with asymptotically flat spacetime. Exploring this second avenue, checking the properties of solutions for a fixed $\mu(\phi)$ is an interesting prospect. Still, we believe that its investigation lies beyond the scope of this work and should be addressed in a future study.
@@
where $G$ is the gravitational constant. From Eqs. (127) and (128), for $n\geq 3/2$, the quantity r(1h)r(1-h) increases at large distances. In this case, we do not have compact SSS objects. For $n$ in the range
@@
the metric component behaves as Eq. (126) and hence $M(r)$ approaches a constant value $c_{1}/(2G)$. This is the case in which nonsingular SSS objects satisfy the condition for compactness.
@@
where the term proportional to $r^{-{\frac{2}{2n-1}}}$ dominates over $r^{-4}$ for $3/4<n<3/2$. As we will see later in Fig. 5, the allowed values of $n$ that are consistent with no-ghost conditions are indeed larger than $3/4$.
@@
In Fig. 2, we plot the ADM mass (129) for seven different values of $n$, with $N_{0}=0.2$. The analytic estimation (126) of $h(r)$ shows that $M$ asymptotically approaches constants for $1/2<n<3/2$, whose property can be confirmed in Fig. 2. For $n=3/2$, the ADM mass exhibits the logarithmic growth $M\propto\ln r$ due to the property of Eq. (127). For $n>3/2$, Eq. (128) gives the analytic dependence $M\propto r^{\frac{2n-3}{2n-1}}$. In Fig. 2, we can confirm this power-law growth of $M$ for $n>3/2$ at large distances. Thus, so long as $1/2<n<3/2$, the electric SSS configuration without the apparent horizon can be interpreted as a compact object.
@@
Figure 5: The allowed region of parameter space for electric compact objects, represented by the light-blue shaded area. The lower boundary corresponds to the values of $n_{\rm min}$, below which the no-ghost condition is violated for a given $N_{0}$ in the range $0<N_{0}<1$. The upper boundary, represented by $n=3/2$, marks the limit beyond which the solutions no longer describe compact objects. Thus, the allowed range for $n$ is $n_{\rm min}\leq n<3/2$. The curve of $n_{\rm min}(N_{0})$ is obtained by interpolating over a set of 99 numerically computed values of $n_{\rm min}$ for $N_{0}$ in the range $0.01\leq N_{0}\leq 0.99$.
@@
The metric component $f=Nh$ differs from $h$ by the factor $N$, whose difference is largest at $r=0$ by the factor $N_{0}=0.2$. Since $N$ grows from $N_{0}$ to 1 for increasing $r$, $f$ approaches $h$ at large distances and they finally converge to the asymptotic values $f\to 1$ and $h\to 1$. The analytic estimation (124) shows that $h(r)$ has the large-distance behavior $h(r)=1+c_{1}/r+c_{2}/r^{2}$ for $n=1$. From Fig. 3, we find that $h(r)<1$ and hence $c_{1}$ is negative. As we see in Fig. 2 for $n=1$, the ADM mass $M$ approaches a constant positive value $c_{1}/(2G)$ at spatial infinity.
@@
Since we are now considering the case $1/2<n<3/2$ with $F>0$, the no-ghost condition for electric compact objects is equivalent to $\mu>0$, i.e.,
$\hskip-25.6073pt\mu$ $=$ $\frac{2^{n-1}M_{\rm Pl}^{2}[2N(1-h-rh^{\prime})-rhN^{\prime}]}{(2n-1)r^{2}N}$ (139)
$\times\left[\frac{q_{E}^{2}(2n-1)^{2}N^{2}}{M_{\rm Pl}^{4}n^{2}\{2(rh^{\prime}+h-1)N+rhN^{\prime}\}^{2}}\right]^{n}>0\,.$
For given values of $N_{0}$, we study how the no-ghost condition $\mu>0$ depends on the parameter $n$ within the allowed range $1/2<n<3/2$. Our numerical analysis reveals the existence of a minimum value of $n$ for each $N_{0}$, denoted as $n_{\rm min}$, above which $\mu>0$ for all $r$. Conversely, if $n<n_{\rm min}$, $\mu$ becomes negative at some finite distance $r$, so these values of $n$ are excluded.
In Fig. 4, we plot $\mu$ versus $r/r_{0}$ with the choice $N_{0}=0.2$ for five different values of $n$. In this case, if $n<0.887$, the coupling enters the region $\mu<0$ for some values of $r$. For $n>n_{\rm min}=0.887$, we observe in Fig. 4 that $\mu>0$ at any distance $r$. For $N_{0}$ in the range $0<N_{0}<1$, we numerically obtain the minimum values of $n$ and plot $n_{\rm min}$ versus $N_{0}$ in Fig. 5. We find that $n_{\rm min}$ mildly increases as a function of $N_{0}$ and approaches 1 for $N_{0}\to 1$. Note that $n_{\rm min}$ is always larger than $1/2$ for any values of $N_{0}$. So long as $n$ is in the range
@@ V.2 Magnetic case
[…]
$\frac{M_{\rm Pl}^{2}hN^{\prime}}{2rN}+\mu(\phi)\left(-\frac{q_{M}^{2}}{2r^{4}}\right)^{n}=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,.$ (143)
If $n$ is an integer, the second term in Eq. (143) reduces to $q_{M}^{2}/(2r^{4})]^{n}$ for even $n$ and $q_{M}^{2}/(2r^{4})]^{n}$ for odd $n$. If $n$ is not an integer, we should think of $F^{n}$ in the Lagrangian as $(F^{2})^{m}$, i.e., $n=2m$, so that Eq. (143) reduces to
@@
For both integer and non-integer values of $n$, the coupling $\mu(\phi)$ is known in terms of $h$, NN, and their $r$ derivatives.
@@
The integrated solution to Eq. (146) can be obtained by replacing the power $n$ to n/(2n1)n/(2n-1) in Eq. (111), so that
@@
Figure 7: This plot depicts the allowed region of parameter space, indicated by the light-blue shaded area. The upper boundary corresponds to the value of $n_{\rm max}$, beyond which the no-ghost condition is violated for a given $N_{0}$ within the allowed range $0<N_{0}<1$. The lower boundary, at $n=3/4$, defines the limit below which the solutions no longer describe a compact object. Consequently, the permissible range for $n$ is $3/4<n\leq n_{\rm max}$.
@@
The metric component $h$ behaves as Eq. (118) around $r=0$, where $n$ should be replaced with n/(2n1)n/(2n-1). At large distances, we require that the ADM mass $M$ is converging, whose condition corresponds to $n<3/2$ for the electric configuration. Using the electric-magnetic duality, the condition for compactness of magnetic SSS objects is now given by $n>3/4$. For the choice (120), the large-distance solutions to $h(r)$ are given by the dual of Eq. (124), that is
@@ VI.1 Electric case
around $r=0$. The reason why the leading-order terms of $c_{r,\delta\phi}^{2}$ and $c_{\Omega,V}^{2}$ are negative is attributed to the property $\kappa\propto F^{-1}$. Indeed, we can obtain the value 1/3-1/3 by substituting the relation $\kappa\propto F^{-1}$ into Eqs. (158) and (159). Even though the dependence of $\kappa$ on $F$ is different for the distance $r$ away from the origin, the Laplacian instability around $r=0$ is sufficient to exclude the above horizonless solution as a stable SSS configuration.
@@ VI.2 Magnetic case
whose leading-order term is negative. Hence the nonsingular SSS object without the horizon is excluded by angular Laplacian instability at large distances. We note that, even without imposing the condition $\tilde{N}_{0}=1$ for asymptotic flatness, the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is 1/2-1/2. The above results show that, as $r$ increases from 0, $c_{\Omega,\delta A}^{2}$ enters the regime $c_{\Omega,\delta A}^{2}<0$ to approach the asymptotic value 1/2-1/2. Indeed, we have numerically confirmed this property [^7] This behavior of $c_{\Omega,\delta A}^{2}$ is similar to what happens for hairy BHs present in cubic vector Galileon theories [64]. by choosing $N(r)$ of the form (120).
@@ VII Conclusions
In Sec. IV, we applied the linear stability conditions to NED with a k-essence scalar field. For nonsingular electric SSS objects, we showed that the even-parity vector-field perturbation $V$ is subject to angular Laplacian instability around $r=0$ due to the negative leading term of $c_{\Omega,V}^{2}$. For the nonsingular magnetic configuration, the same angular Laplacian instability arises for the odd-parity vector-field perturbation $\delta A$. Since the fields $V$ and $\delta A$ are coupled to the even-parity gravitational perturbation $\psi$, the background metric regular at $r=0$ cannot remain in a steady state. This means that, in theories given by the Lagrangian $\tilde{\cal L}(F)+K(\phi,X)$, the nonsingular SSS objects (including regular BHs) do not exist as stable configurations.